OpenStudy (omnomnom):

The January 14, 2002 issue of TIME magazine reported that Wal-Mart’s 3rd quarter revenues for 2001 were 600% more than its No. 2 competitor, K-Mart. What were K-Mart’s revenues during this period if Wal-Mart’s revenues were $48 billion?

OpenStudy (anonymous):

@Hero @TuringTest @dpaInc

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@ash2326 @UnkleRhaukus @phi

OpenStudy (anonymous):

@radar @Agent_Sniffles @AccessDenied @karatechopper

OpenStudy (anonymous):

This is actually my question but i couldnt leave my post since i need it to review

hero (hero):

K-Mart is Walmart's #2 competitor? That has to be a mistake.

OpenStudy (anonymous):

its just a math problem hero!

OpenStudy (anonymous):

in reality its probably costco or target

OpenStudy (phi):

you need to find 600% more than 48

OpenStudy (anonymous):

how do i go about doin that

OpenStudy (phi):

The first step is find 600% of 48 if the problem were 10% of 48 can you do it? it is the same idea: 600% times 48 change 600% to a fraction: 600/100 simplify 6 multiply 6*48

OpenStudy (phi):

step 2. 600% more than 48 means add that amount to 48 so add the answer from step 1 to 48 to get the final answer

OpenStudy (anonymous):

oh its not 6/48?

OpenStudy (phi):

No. one way to remember is use simple numbers that you understand you probably know 50% is the same as 1/2 , right (I want 50% means I want half) 50% of 48 is 1/2 of 48 or 24 so if we do the simple procedure: make 50% a fraction over 100: 50/100 simplify: 1/2 multiply 1/2*48= 24 so multiplying is the right thing to do, we get the correct answer

OpenStudy (anonymous):

6 times 48 is 288 thats such a big number

OpenStudy (anonymous):

what is the next step

OpenStudy (phi):

too bad it isn't your bank account. now you want 600% more than 48 so the answer is 288 more than 48 this is just a way to say 288+48

OpenStudy (anonymous):

i have the answer btw i just need to know how to do it and its not 288

OpenStudy (anonymous):

it comes out to be 336 but thats not the answer....

OpenStudy (phi):

that's because I read it backwards wal-mart is 48 and they want k-mart 600% of Kmart + Kmart = 48 now we have to find Kmart can you do this?

OpenStudy (anonymous):

uhhh yea if you weren't so vague =/

OpenStudy (phi):

change 600% to 600/100 and then to 6 6*K + K = 48 add 6K plus 1K to get 7K 7K= 48 now divide by 7 to get K (Kmart)

OpenStudy (anonymous):

where did the 1k come from?

OpenStudy (phi):

I was thinking that 600% more than K means K plus 600% of K

OpenStudy (anonymous):

you're soooooooooooooooooo confusing me right now

OpenStudy (anonymous):

the answer is 6.86 billion i have no clue how that is

OpenStudy (phi):

Let's say that Walmarts make 10% more than K it would make K and then 10% more now it's 600% more, so K + 600% of K K+6K = 48 7K= 48 K= 48/7

OpenStudy (anonymous):

wut is this k u speak of

OpenStudy (phi):

K is short for how much Kmart makes. It's easier to type K than all of that

OpenStudy (anonymous):

um ok i still dont understand where the 7 came form

OpenStudy (phi):

Do you know the answer to this: Your friend make $100 you made 10% more than your friend how much did you make?

OpenStudy (anonymous):

110

OpenStudy (phi):

now if we write that as an equation: you = friend + 10% of friend (where you means how much you made, and friend means how much your friend made)

OpenStudy (phi):

You see it works: you = 100 + 10% of 100 you = 100 + 10/100 * 100 you = 100+ 10 you=110

OpenStudy (anonymous):

man see u just lose me when u pull that up =(

OpenStudy (phi):

now say you make 600% more than your friend. we use the same idea, but different numbers you = friend + 600% of friend you = friend + 6 * friend (we change 600% to 600/100 or 6) now in algebra we can add 1 friend + 6 friend and get 7 friends (this is money not friends) you = 7 friend

OpenStudy (anonymous):

can u just cut the friend thing u're losing me

OpenStudy (phi):

do you know if you have x+6x you can simplify to 7x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

with a 1

OpenStudy (phi):

so your question is how did we get to x+6x where x is how much kmart makes in one quarter The answer is we change the words 600% more than K-Mart. into "math" I tried to show that 10% more than your friend means : what your friend made plus 10% of what your friend made here if x is how much K-mart made 600% more means what K-mart made plus 600% of what K-mart made

hero (hero):

Wow, speechless

OpenStudy (anonymous):

oh ok i kinda ge tit now

OpenStudy (omnomnom):

O.O

OpenStudy (phi):

Think about it, and practice..